OpenStudy (anonymous):

write it in radical form (6v)^1.5

OpenStudy (mathstudent55):

Remember that 1.5 = 3/2 Also, here is an important rule: \(\large a^{\frac{m}{n}} = \sqrt[n]{a^m} \)

OpenStudy (anonymous):

okay thanks

OpenStudy (mathstudent55):

You're welcome.